# Large Language Model Based Multi-Objective Optimization for Integrated Sensing and Communications in UAV Networks

### v1

@@ V Conclusions
We studied a multi-UAV-enabled ISAC system to maximize the total network utility and minimize the localization CRBs. We leveraged the benefits of integrating LLMs with MOEAs by proposing an LLM-enabled decomposition-based MOEA (LEDMA) to solve the formulated MOP, in which the original MOP was decomposed into a series of sub-problems, and the LLMs are instructed as search operators with MOP-specifically designed prompts to solve sub-problems simultaneously. Numerical results demonstrated that our proposed algorithm outperformed the baseline MOEAs in terms of acquiring Pareto fronts and convergence. Our results also demonstrated the significant potential of LLMs to solve optimization problems in wireless communications. In our future work, we will extend our work into mobile user scenarios and explore to what extent the scalability of the network topology can affect the effectiveness of the proposed algorithm.

### v2

@@ V Conclusions
We studied a multi-UAV-enabled ISAC system to maximize the total network utility and minimize the localization CRBs. We leveraged the benefits of integrating LLMs with MOEAs by proposing an LLM-enabled decomposition-based MOEA (LEDMA) to solve the formulated MOP, in which the original MOP was decomposed into a series of sub-problems, and the LLMs are instructed as search operators with MOP-specifically designed prompts to solve sub-problems simultaneously. Numerical results demonstrated our proposed algorithm outperformed the baseline MOEAs in terms of acquiring Pareto fronts and convergence. Our results also demonstrated the significant potential of LLMs to solve optimization problems in wireless communications. In our future work, we will extend our work into mobile user scenarios and explore to what extent the scalability of the network topology can affect the effectiveness of the proposed algorithm.